# Discrete Curvature on Graphs

### v1

@@ 1 Introduction
The reverse isopermietric problem is inspired from the classical isoperimetric problem which has been thoroughly studied in the past and generalized in a variety of contexts (see [1]). The classical problem concerns about finding the body with the largest volume for a given surface area. In $\mathbb{R}^{n}$ it is well known that the solution is an $n$-dimensional ball. The reverse isoperimetric problem addresses the opposite case. That is, it searches for the body that minimizes the volume for a given surface area. This problem has a trivial solution as, for a given surface area, the $n$-dimensional volume can be arbitrary small (the flat pancake problem [2]).
To avoid this degeneration, in this paper we restrict the reverse isoperimetric problem to convex bodies with uniform curvature constraints. Our focus will be on bodies lying in the hyperbolic plane $\mathbb{H}^{2}$, which we assume has constant curvature of -$1$. In particular, we look at convex bodies $K\subset\mathbb{H}^{2}$ whose boundary curvature satisfies $1/\lambda\leq\kappa\leq\lambda$, in a weak sense (see below). We call these thick $\lambda$-concave bodies. Recall that a convex body is a closed convex set with non-empty interior. This restricts the body from being too flat or too curved.
@@ 1.1 Some reverse isoperimetric results
In this paper, we focus on $\lambda$-concave and $\frac{1}{\lambda}$-convex bodies. Throughout, we assume that $\lambda>1$ and that the normal to the boundary $\partial K$ of a convex body $K$ points in the direction of convexity.
Definition 1 ($\lambda$-convex and $\lambda$-concave bodies [3]).
A convex body $K\subset\mathbb{H}^{2}$ is $\lambda$-convex if for each $x\in\partial K$ there exists a neighborhood $U_{x}\subset\mathbb{H}^{2}$ and a closed curve of constant curvature $\partial S_{\lambda}$ (see Section 2.1) such that
$U_{x}\cap\partial K\subset S_{\lambda}$
@@
$U_{x}\cap\partial S_{\frac{1}{\lambda}}\subset K$
Equivalently, if $\partial K$ is $C^{2}$-smooth then $K$ is $\lambda$-convex if for all $x\in\partial K$ the geodesic curvature $k(x)\geq\lambda$, and $\lambda$-concave if for all $x\in\partial K$ the geodesic curvature $0\leq k(x)\leq\lambda$.
@@
  • The Reverse Isoperimetric Inequality for $\lambda$-concave Bodies ([4], see also [5, 6]). Let $n\geq 2$. Let $K\subset\mathbb{R}^{n}$ be a $\lambda$-concave body, and let $S\subset\mathbb{R}^{n}$ be a $\lambda$-sausage, i.e., the convex hull of two balls of radius $1/\lambda$. If $|\partial K|=|\partial S|$, then $|K|\geq|S|$, and equality holds if and only if $K$ is a $\lambda$-sausage.
@@
  • Conjecture: The Reverse Isoperimetric Inequality for $\lambda$-convex bodies in $\mathbb{R}^{n}$. Let $n\geq 2$. Let $K\subset\mathbb{R}^{n}$ be a $\lambda$-convex body, and let $L\subset\mathbb{R}^{n}$ be a $\lambda$-convex lens, i.e., the intersection of two balls of radius $1/\lambda$. If $|\partial K|=|\partial L|$, then $|K|\geq|L|$, and equality holds if and only if $K$ is a $\lambda$-convex lens.
@@
For $n=2$ the conjecture was proved by [7] (see [8] for alternative proof). For $n=3$ the conjecture was proved in [3]. In this paper, we also address the Blaschke’s Rolling Theorem for thick $\lambda$-concave bodies. The statement of the classical result is as follows:
  • Theorem: Blaschke’s Rolling Theorem [9]. Let $\gamma$ be a closed, smooth, convex curve in $\mathbb{R}^{2}$ such that for all $x\in\gamma,\,\kappa(x)\leq\lambda$. Then, a ball of curvature $\lambda$ rolls freely inside.
This result was extended in $\mathbb{R}^{n}$ by [10, 11], and later on in Riemannian manifolds of bounded curvature by [12, 13, 14, 15]. Nevertheless, this theorem does not hold in general for $\lambda$-concave domains in $\mathbb{H}^{2}$, and hence was never studied in literature before. We fix this gap by providing a Blaschke’s Rolling Theorem for $\lambda$-concave bodies under the thickness condition (Theorem C). We also show that this result is optimal.
@@ 1.2 Reverse isoperimetric inequality in $\mathbb{H}^{2}$
Definition 2 (Thick $\lambda$-sausage).
@@
In this paper we provide a proof of the theorems stated below. Note that given $K\subset\mathbb{H}^{2}$, $|\partial K|$ refers to the length of the boundary $\partial K$ and $|K|$ to the area of $K$.
Theorem A.
(Reverse isoperimetric inequality for thick $\lambda$-concave bodies in $\mathbb{H}^{2}$) Let $K\subset\mathbb{H}^{2}$ be a thick $\lambda$-concave body and let $S\subset\mathbb{H}^{2}$ be a thick $\lambda$-sausage, as in Definition (2). If $|\partial K|=|\partial S|$, then $|K|\geq|S|$. Moreover, equality holds if and only if $K$ is a thick $\lambda$-sausage.
Theorem B.
(Lower bound for reverse isoperimetric inequality for thick $\lambda$-concave bodies in $\mathbb{H}^{2}$) Let $K\subset\mathbb{H}^{2}$ be a thick $\lambda$-concave body. Then,
@@
Remark: In fact, if $K\subset\mathbb{H}^{2}(-c^{2})$ is a thick $\lambda$-concave body, then,
@@
And, as $c\rightarrow 0$, the inequality approaches the classical Euclidean result in $\mathbb{R}^{2}$ proved in [4] (this inequality complements the analogous results in $\mathbb{S}^{2}$ [16])
Theorem C.
(Blaschke’s rolling theorem for $\lambda$- concave bodies) Let $K\subset\mathbb{H}^{2}$ is a thick $\lambda$-concave body. Then the ball of curvature $\lambda$ can roll freely in $K$. Moreover, for every $\lambda^{\prime}>\lambda$, there exists a closed convex curve with $1/\lambda^{\prime}\leq k\leq\lambda$ such that a disk of radius $R=\coth^{-1}(\lambda)$ can not roll freely in the domain bounded by the curve.
@@ 1.3 Structure of the paper
In Section 2, we provide some background information on the curvature of convex bodies in the hyperbolic plane. In Section 3 we give the proof of Theorem A, which is divided into three steps for clarity. The proof follows a similar approach than in the Euclidean case, where the structure of the inner parallel bodies is analyzed, but with the additional difficulty than in the hyperbolic plane the inner parallel body of a convex body is not necessarily convex. The proof of Theorem B and Theorem C will mostly follow from Theorem A.
@@ 2.1 Curves of constant curvature in $\mathbb{H}^{2}$
In the Hyperbolic Space of curvature -$1$ there are multiple types of curves of constant geodesic curvature (see Figure 3). This section provides the definition and properties of such curves (see e.g [17] or [3]).
  • Circles: Circles of radius $R$ have geodesic curvature $\lambda=\lambda=\coth(R)>1$. Both in the Poincaré Disk Model and in the Upper Half-Plane Model, hyperbolic circles are Euclidean circles, even though, in general, the hyperbolic center does not coincide with the Euclidean center of the disk.
  • Geodesics: The geodesic curvature of geodesics is 0. In the Poincaré Disk Model, geodesics appear as arcs of circles that are orthogonal to the boundary of the disk or as straight lines passing through the center. In the Upper Half-Plane Model, geodesics are represented as either semicircles orthogonal to the boundary (the real line) or vertical lines perpendicular to the boundary.
@@
Figure 3: Hyperbolic (green), Geodesics (orange), Horocycles (blue) and Hypercircles (pink)
@@ 2.2 Some useful lemmas
In the Poincaré Disk Model, the geodesic curvature of a curve at (0,0) is $\frac{1}{2}$ of its Euclidean curvature at that point.
@@
The Poincaré metric in the unit disk $x^{2}+y^{2}<1$ is $ds^{2}=\frac{4\left(dx^{2}+dy^{2}\right)}{\left(1-x^{2}-y^{2}\right)^{2}}$. To calculate the geodesic curvature of a curve through (0,0), we use the parallel transport to quantify how much does the curve curve with respect to the geodesic, which in this case is the diameter of the circle. Near the center, its arclength parametrization is approximately $(\frac{t}{2},0)$ for $t\rightarrow 0$. Hence, the parallel transport along this geodesic for small distances near center will be Euclidean and the geodesic curvature at that point will be just $\frac{1}{2}$ of its Euclidean curvature at that point. ∎
@@
Let $\beta$ be the angle between the hypercicle and the boundary of the of the unit disk in the Poincaré Disk Model, as seen in Figure 4. Then, the curvature $\lambda$ of an hypercircle $\Gamma\subset\mathbb{H}^{2}$ is $\lambda=\cos(\beta)$.
@@
Let $\Gamma$ be a hypercircle that passes through the center of the disk (note that we can map any hypercircle to the center of this disk with a Möbius transformation). We can then consider the construction shown in Figure 4:
Figure 4: Construction of the proof of Lemma 2.2. We consider the triangle joining the center of the unit disk, the center of the hypercirlce and their point of intersection. Using properties of the tangent lines at the point of intersection, we can relate the radius R to the curvature at (0,0) and $\beta$.
First, notice that $\beta=\alpha$. This is because the radius at the point of tangency is perpendicular to the tangent (red lines). Therefore, we get that $\delta=\pi-\alpha$, which leads to $\delta+\beta=\pi$ and $\beta=\alpha$. Now, let us construct the blue triangle as shown in Figure 4. Using some trigonometric identities we get that
@@
The hypercircle $\Gamma$ at (0,0) has Euclidean curvature $\frac{1}{R}=2\cos(\alpha)$. Therefore, from Lemma 2.1, the hyperbolic curvature of $\Gamma$ at (0,0) is $\cos(\alpha)=(\cos(\beta))$. Since (0,0) is not a special point, every point of $\gamma$ has hyperbolic curvature $\cos(\beta)$. ∎
@@
We can parameterize these family of arcs as:
@@
for $A>0$. Then, $R$ can be obtained by calculating the hyperbolic length of $\sigma(t)$ from $\beta\leq t\leq\frac{\pi}{2}$,
$R=\int_{\beta}^{\frac{\pi}{2}}\frac{|\sigma^{{}^{\prime}}(t)|}{\text{Im}(\sigma(t))}\text{dt}=\int_{\beta}^{\frac{\pi}{2}}\frac{1}{\sin(t)}\text{dt}=\log(\tan(\frac{\beta}{2})).$
@@
$\tanh(R)=\tanh(\log(\tan(\frac{\beta}{2})))=-\cos(\beta).$
The negative sign comes from the direction of the normal component. Therefore, $\tanh(R)=|\cos(\beta)|$ as required. Note that when $\Gamma\subset\mathbb{H}^{2}(-c^{2})$ then the curvature $\lambda$ of $\Gamma$ is normalized to $\lambda=c\cdot\tanh(cR)$. ∎
@@
Let $\Gamma$ be the hypercircle a distance $R$ from the geodesic joining the center of two balls of radius $R$ and curvature $\lambda$. Then, the curvature of the hypercircle is $\frac{1}{\lambda}$.
@@
The curvature of a circle of radius $R$ is $\coth(R)$. From Lemma 2.3, the curvature of the hypercircle is $\tanh(R)=\tanh(\coth^{-1}(\lambda))=\frac{1}{\lambda}$, as required.
3 Proof of Theorems A, B and C
The proof of Theorem A is divided into 3 steps. The strategy is as follows:
@@ 3 Proof of Theorems A, B and C
    Given a thick $\lambda$-concave body we construct a body $H$ named the Thick Convex Hull of $K$ such that $H\subset K$. We do that by analyzing the structure of the cut locus $T$ of $K$ (subsection 3.1).
@@
    We then define the inner parallel bodies of both $H$ and $K$ (subsection 3.2).
  3. 3.
@@
    We use the classical Steiner’s formula to relate the area between these two bodies (subsection 3.3) and their inner parallel bodies.
3.1 Construction of the Thick Convex Hull of $K$
The first step is to construct a body $H$ (named the Thick Convex Hull of $K$), such that $H\subset K$. We will do that by analyzing the structure of the cut locus $T$ of $\partial K$. The following definitions are needed for the analysis (see [18]):
Definition 3 (Cut Locus).
Let $K$ be a connected domain with picewise smooth boundary $\partial K$. The cut locus $T$ of $K$ with respect to its boundary $\partial K$ is defined as the closure of the set of all points $x\in K$ such that the restriction of the distance function $d(x,\partial K)$ attains its global minimum at two or more points of $\partial K$ or, in other words, for which there exist two or more points in $\partial K$ minimizing the distance to $x$ [18].
Definition 4 (Focal Point).
Let $K$ be a domain with smooth boundary $\partial K$. For each point $x$ belonging to a cut locus $T$, consider the subset $X\subset\partial K$ where the minimal distance to $x$ is attained. We say that $x$ is a focal point if these equivalent statements are satisfied [18]:
  • $X$ is connected
  • For any sufficiently small neighborhood $U\ni x$, the complement $U\backslash T$ is connected
If $x\in T$ is not a focal point, then we say that is a cut point.
An illustration of the cut locus and focal points of $\partial K$ for $K\subset\mathbb{H}^{2}$ is shown in Figure 6.
Figure 6: Cut locus of $\partial K$
Definition 5 (Strong Deformation Retract).
A subspace $A$ of a topological space $X$ is a strong deformation retract of $X$ if there is a homotopy $F:X\times I\to X$such that:
$F(x,0)=x\quad\forall x\in X$
$F(a,t)=a\quad\forall a\in A,t\in I$
$F(x,1)\in A\quad\forall x\in X$
Lemma 3.1 (Structure of the cut locus away from focal points [19]).
Let $K\in\mathbb{H}^{2}$ be a convex domain with cut locus $T$ and let $P\in T$ be a cut point that is not a focal point with $R=\text{d}(P,\partial K)$. Then there is a finite number of $k\geq 2$ of minimizing geodesics from $P$ to $\partial K$, and
  • Case 1. If $k=2$, then there is a neighborhood $U$ of $P$ so that $T\cap U$ is a $C^{1}$ curve and the tangent to $T$ at $P$ bisects the angle between the two minimizing geodesics from $P$ to $\partial K$.
  • Case 2. If $k\geq 3$, then the $k$ geodesic segments from $P$ to $\partial K$ split the disk $B_{R}(P)$ into $k$ sectors $S_{1},\dots,S_{k}$. There is a small open disk $U$ around $P$ so that in each sector $S_{i}$ the set $T\cap U\cap S_{i}$ is a $C^{1}$ curve ending at $P$ and the tangent to this curve at $P$ is the angle bisector of the two sides of the sector $S_{i}$ at $P$ (see Figure 7).
The proof of Lemma 3.1 can be found in [19]. Below there is a diagram showcasing the statement of the Lemma.
Figure 7: Diagram of Lemma 3.1. The orange line is the cut locus $T$ of $\partial K$ and the slashed blue lines are the 3 geodesics from $P$ to $\partial K$.
Note that from Lemma 3.1, any vertex of $T$ which is not a focal point must be of valance at least 3.
Lemma 3.2.
Let $K$ be a convex body in $\mathbb{H}^{2}$. The cut locus $T$ is a strong deformation retract of $K$ [19].
This retraction can be obtained by considering each point $y^{\prime}\in K\backslash T$ and the corresponding point $y\in\partial K$ that minimizes the distance d($y,y^{\prime}$). We can then move $y$ towards $T$ along the geodesic $[y,y^{\prime}]$. A consequence of Lemma 3.2 is that $K$ and $T$ are homotopy equivalent. In particular, if $K$ is simply connected, then $T$ must also be simply connected and, hence, a tree.
Lemma 3.3.
Let $K$ be a $\lambda$-concave body different from a disk. Then, there exists at least two disks of radius $\coth^{-1}(\lambda)$ contained in $K$.
Proof.
We may assume that $T$ is a tree embedded in $K$ with a finite number of edges. From Lemma 3.1 any vertex of degree 1 must be a focal point. Hence, it contains at least two focal points which corresponds to the leaves of $T$ (except when $K$ is a disk, in which case, it only contains one focal point). In addition, from Definition 4 if $x$ is a focal point, then there exists a $z\in\partial K$ such that the osculating circle at $z$ is centered at $x$. Since $\partial K$ has curvature at most $\lambda$, the radius of the osculating circle is at least $R=1/c\cdot\coth^{-1}(\lambda/c)$. Hence, the distance from $x$ to $\partial K$ is at least $R$, which proves that in each leave of $T$ there is an osculating circle $B_{R}$, of radius $R$, such that $B_{R}\subset K$. ∎
Lemma 3.4.
Let $K\subset\mathbb{H}^{2}$ be $\lambda$-concave body different than a disk. Then,
$|K|>4\pi\left(\frac{1}{\sqrt{1-\frac{1}{\lambda^{2}}}}-1\right)$
Proof.
From Lemma 3.3, we have that a $K$ contains at least 2 disks of radius $R=\coth^{-1}(\lambda)$. We can then use the formula for the area of a disk in the hyperbolic space to get the desired inequality. ∎
Definition 6 (Thick Convex Hull of T).
Let $T$ be the cut locus of a thick $\lambda$-concave body $K\subset\mathbb{H}^{2}(-c^{2})$ and let $V(T)$ be the set of vertices of $T$. For all $i\in V(T)$ consider the ball $B_{i}$ centered at the vertex $i$ and with curvature $\lambda_{i}$, where $\lambda_{i}$ is the minimum curvature such that $B_{{}_{i}}\subset K$ (note that $\lambda_{i}\leq\lambda\,\,\forall i$ ). The Thick Convex Hull of $T$ of curvature $\lambda$ is defined as
$\text{ThickConv }(T,\lambda):=\bigcup_{x,y\in W}[x,y]_{\lambda}$
where $W$ is the convex hull of $\bigcup B_{i}$ and $[x,y]_{\lambda}$ is the region bounded by two hypercircles of curvature $1/\lambda$ joining $x$ and $y$.
Theorem 1.
[…]
The proof of Theorem 1 can be found in [20]. From this Theorem we get that ThickConv ($\text{ThickConv }(T,\lambda)\subset K$. Figure 8 shows ThickConv ($\text{ThickConv }(T,\lambda)$ (where $\lambda$ is the curvature of the circles centered at the vertices of $T$) and its respective convex hull $W$.
Figure 8: Example of a Thick Convex Hull of $T$ and convex hull (region bounded by blue lines and orange lines, respectively)
Remark: Notice that it is not possible for a ball $B_{\lambda}$ (ball of curvature $\lambda$) centered at vertex of $T$ to be fully contained in the interior of ThickConv ($\text{ThickConv }(T,\lambda)$, like the situation exemplified in the Figure 9 . This is because from Definition 4, if a vertex $v$ is a focal point, there exists a subset $X\subset\partial K$, for $K=\text{ThickConv }(T,\lambda)$, such that the minimal distance to $v$ is achieved for all $x\in X$.
Figure 9: Non-possible construction of the Thick Convex Hull

### v2

# Reverse Isoperimetric Properties of Thick $\lambda$-Concave Bodies in the Hyperbolic Plane

@@ Abstract
In this paper we address the reverse isoperimetric inequality for convex bodies with uniform curvature constraints in the hyperbolic plane $\mathbb{H}^{2}$. We prove that the thick $\lambda$-sausage body, that is, the convex domain bounded by two equal circular arcs of curvature $\lambda$ and two equal arcs of hypercircle of curvature $1/\lambda$, is the unique minimizer of area among all bodies $K\subset\mathbb{H}^{2}$ with a given length and with curvature of $\partial K$ satisfying $1/\lambda\leq\kappa\leq\lambda$ (in a weak sense). We call this class of bodies thick $\lambda$-concave bodies, in analogy to the Euclidean case where a body is $\lambda$-concave if $0\leq\kappa\leq\lambda$. The main difficulty in the hyperbolic setting is that the inner parallel bodies of a convex body are not necessarily convex. To overcome this difficulty, we introduce an extra assumption of thickness $\kappa\geq 1/\lambda$.
Keywords: reverse isoperimetric inequality, $\lambda$-concavity, $\lambda$-convexity, hyperbolic plane
@@ 1 Introduction
The reverse isopermietric problem is inspired from the classical isoperimetric problem which has been thoroughly studied in the past and generalized in a variety of contexts (see [1]). The classical problem concerns with finding the body with the largest volume for a given surface area. In $\mathbb{R}^{n}$ it is well known that the solution is an $n$-dimensional ball. The reverse isoperimetric problem addresses the opposite case. That is, it searches for the body that minimizes the volume for a given surface area. This problem has a trivial solution as, for a given surface area, the $n$-dimensional volume can be arbitrary small (the flat pancake problem [2]).
To avoid this degeneration, in this paper we restrict the reverse isoperimetric problem to convex bodies with uniform curvature constraints. Our focus will be on convex bodies lying in the hyperbolic plane $\mathbb{H}^{2}$, which we assume has constant curvature of 1-1. Recall that a set is called convex if the geodesic segment between any two points in it, lies inside. A convex body is a compact convex set with non-empty interior. In particular, we look at convex bodies $K\subset\mathbb{H}^{2}$ whose boundary curvature satisfies $1/\lambda\leq\kappa\leq\lambda$, in a weak sense (see below). We call these thick $\lambda$-concave bodies. This curvature constraint restricts the body from being too flat or too curved, ruling out trivial solutions such as the “flat pancake shape”.
@@ 1.1 Some reverse isoperimetric results
In this paper, we focus on $\lambda$-concave and $1/\lambda$-convex bodies. Throughout, we assume that $\lambda>1$ and that the normal vector to the boundary $\partial K$ of a convex body $K$ points in the direction of convexity. In addition, given a convex body $K$, $|\partial K|$ refers to the length (or surface area) of the boundary $\partial K$ and $|K|$ to the area (or volume) of $K$.
Definition 1.1 ($\lambda$-convex and $\lambda$-concave bodies in $\mathbb{H}^{2}$ [3]).
A convex body $K\subset\mathbb{H}^{2}$ is $\lambda$-convex if for each $x\in\partial K$ there exists a neighborhood $U_{x}\subset\mathbb{H}^{2}$ and a closed curve $\partial S_{\lambda}$ of constant curvature $\lambda$ passing through $x$ (see Section 2.1) such that
$U_{x}\cap\partial K\subset S_{\lambda}.$
@@
$U_{x}\cap\partial S_{\lambda}\subset K.$
Equivalently, if $\partial K$ is $C^{2}$-smooth, then $K$ is $\lambda$-convex when the geodesic curvature $\kappa_{g}$ satisfies $\kappa_{g}(x)\geq\lambda$ for all $x\in\partial K$, and $\lambda$-concave when the geodesic curvature $\kappa_{g}$ satisfies $0\leq\kappa_{g}(x)\leq\lambda$ for all $x\in\partial K$.
Definition 1.2 (Thick $\lambda$-concave bodies).
A convex body $K\subset\mathbb{H}^{2}$ is thick $\lambda$-concave if the geodesic curvature $\kappa_{g}$ satisfies $1/\lambda\leq\kappa\leq\lambda$ (in a weak sense).
@@
Theorem 1.1 (Reverse Isoperimetric Inequality for $\lambda$-concave Bodies).
([4], see also [5, 6]). Let $n\geq 2$. Let $K\subset\mathbb{R}^{n}$ be a $\lambda$-concave body, and let $S\subset\mathbb{R}^{n}$ be a $\lambda$-sausage, i.e., the convex hull of two balls of radius $1/\lambda$. If $|\partial K|=|\partial S|$, then $|K|\geq|S|$, and equality holds if and only if $K$ is a $\lambda$-sausage.
@@
Conjecture (The Reverse Isoperimetric Inequality for $\lambda$-convex bodies in $\mathbb{R}^{n}$). Let $n\geq 2$. Let $K\subset\mathbb{R}^{n}$ be a $\lambda$-convex body, and let $L\subset\mathbb{R}^{n}$ be a $\lambda$-convex lens, i.e., the intersection of two balls of radius $1/\lambda$. If $|\partial K|=|\partial L|$, then $|K|\geq|L|$, and equality holds if and only if $K$ is a $\lambda$-convex lens.
@@
For $n=2$ the conjecture was proved in [7] (see [8] for an alternative proof). For $n=3$ the conjecture was proved in [3].
@@ 1.2 Reverse isoperimetric inequality in $\mathbb{H}^{2}$
Definition 1.3 (Thick $\lambda$-sausage).
@@
In this paper we provide proofs of the theorems stated below.
Theorem 1.2 (Reverse isoperimetric inequality for thick $\lambda$-concave bodies in $\mathbb{H}^{2}$).
Let $K\subset\mathbb{H}^{2}$ be a thick $\lambda$-concave body and let $S\subset\mathbb{H}^{2}$ be a thick $\lambda$-sausage, as in Definition 1.3. If $|\partial K|=|\partial S|$, then $|K|\geq|S|$. Moreover, equality holds if and only if $K$ is a thick $\lambda$-sausage.
Theorem 1.3 (Lower bound for reverse isoperimetric inequality for thick $\lambda$-concave bodies in $\mathbb{H}^{2}$).
Let $K\subset\mathbb{H}^{2}$ be a thick $\lambda$-concave body. Then,
@@
Remark 1.4.
In fact, if $K\subset\mathbb{H}^{2}(-c^{2})$ is a thick $\lambda$-concave body in the hyperpolic plane of curvature $c^{2}$, $c>0$, then
@@
And, as $c\rightarrow 0$, the inequality approaches the classical Euclidean result in $\mathbb{R}^{2}$ proved in [4]. This inequality also complements the analogous results obtained in $\mathbb{S}^{2}$ [9].
@@ 1.3 Structure of the paper
In Section 2, we provide some background theory on the curvature of convex bodies in the hyperbolic plane. In Section 3, we give the proof of Theorem 1.2 and Theorem 1.3. The proof follows a similar approach as in the Euclidean case, where the structure of the inner parallel bodies is analyzed, but with the additional difficulty that in the hyperbolic plane the inner parallel bodies of a convex body are not necessarily convex.
@@ 2.1 Curves of constant curvature in $\mathbb{H}^{2}$
In the hyperbolic space of curvature -$1$ there are multiple types of curves of constant geodesic curvature (see Figure 3). This section provides the definition and key properties of such curves in the Poincaré Disk model and the Upper Half-Plane model, which are the relevant models for our analysis (see e.g Section 35.2.4 in [10] or Section 2.1 [3]).
  • Circles: Circles of radius $R$ have geodesic curvature $\lambda=\coth(R)>1$. Both in the Poincaré Disk model and in the Upper Half-Plane model, hyperbolic circles are Euclidean circles, even though, in general, the hyperbolic center does not coincide with the Euclidean center of the disk.
  • Geodesics: By definition, the geodesic curvature of geodesics is 0. In the Poincaré Disk model, geodesics appear as arcs of circles that are orthogonal to the boundary of the disk or as straight lines passing through the center. In the Upper Half-Plane model, geodesics are represented as either semicircles orthogonal to the boundary (the real line) or vertical lines perpendicular to the boundary.
@@
Figure 3: Circle (green), Geodesics (orange), Horocycles (blue) and Hypercircles (pink)
@@ 2.2 Some useful lemmas
In the Poincaré Disk model, the geodesic curvature of a curve at the origin is $\frac{1}{2}$ of its Euclidean curvature at that point.
@@
This result is standard, and we will only give a heuristic idea of the proof. To compute the geodesic curvature of a curve at $(0,0)$ in the Poincaré disk, we use parallel transport to quantify how much the curve deviates from a geodesic. We take the geodesic to be the horizontal diameter of the disk. Let $t$ denote the hyperbolic arclength parameter along the curve. Near the origin, the hyperbolic metric is approximately Euclidean up to a scaling factor of 2, so the Euclidean coordinates of the curve are approximately $\left(\frac{t}{2},0\right)$ for $t\to 0$. Since parallel transport along the geodesic near the origin behaves approximately as in Euclidean space, the geodesic curvature at $(0,0)$ is half the Euclidean curvature at that point. ∎
@@
Let $\beta$ be the angle at which the hypercircle $\Gamma$ intersects the boundary of the unit disk in the Poincaré Disk model, as illustrated in Figure 4. Then the curvature $\lambda$ of $\Gamma$ in $\mathbb{H}^{2}$ is given by $\lambda=\cos(\beta).$
@@
Let $\Gamma$ be a hypercircle containing the center of the disk (note that we can map any hypercircle to the center of this disk with a Möbius transformation). We can then consider the construction shown in Figure 4, where the black circle denotes the boundary of the Poincaré Disk and the green circle the hypercircle $\Gamma$.
Let $\delta$ denote the angle between the tangent to the unit circle and the tangent to the hypercircle at their point of intersection (red lines in Figure 4). Since the radius of the circle is perpendicular to its tangent at the boundary point, we have that $\delta=\pi-\alpha$, which leads to $\delta+\beta=\pi$. This implies that $\beta=\alpha$. Using some trigonometric identities, we get that
@@
The hypercircle $\Gamma$ at (0,0) has Euclidean curvature $1/R=2\cos(\alpha)$. Therefore, from Lemma 2.1, the hyperbolic curvature of $\Gamma$ at (0,0) is $\cos(\alpha)=\cos(\beta)$. Since (0,0) is not a special point, every point of $\gamma$ has hyperbolic curvature $\cos(\beta)$. ∎
Figure 4: The proof of Lemma 2.2 is constructed as follows. We consider the triangle formed by the center of the unit disk (black), the center of the hypercircle (green), and their point of intersection. By examining the properties of the tangent lines to the disk and hypercircle at the point of intersection, we can establish a relationship between the radius $R$, the curvature at (0,0) and $\beta$.
@@
We can then parameterize this family of arcs as:
@@
for $A>0$. Then, we can calculate $R$ by computing the hyperbolic length of $\sigma(t)$ from $\beta\leq t\leq\frac{\pi}{2}$,
$R=\int_{\beta}^{\frac{\pi}{2}}\frac{|\sigma^{{}^{\prime}}(t)|}{\text{Im}(\sigma(t))}\text{dt}=\int_{\beta}^{\frac{\pi}{2}}\frac{1}{\sin(t)}\text{dt}=-\log\left(\tan\left(\frac{\beta}{2}\right)\right).$
@@
$\tanh(R)=\frac{1-e^{-2R}}{1+e^{-2R}}=\frac{1-\tan^{2}\left(\frac{\beta}{2}\right)}{1+\tan^{2}\left(\frac{\beta}{2}\right)}=\cos(\beta).$
Note that when $\Gamma\subset\mathbb{H}^{2}(-c^{2})$ then the curvature $\lambda$ of $\Gamma$ is normalized to $\lambda=c\cdot\tanh(cR)$
@@
Let $\Gamma$ be a hypercircle that lies at a distance $R$ from a geodesic. Suppose this geodesic connects the centers of two disks, each of radius $R$ and curvature $\lambda$ (see Figure 2). Then the geodesic curvature of $\Gamma$ is equal to $1/\lambda$.
@@
The curvature of a circle of radius $R$ is $\coth(R)$. From Lemma 2.3, the curvature of the hypercircle is $\tanh(R)=\tanh(\coth^{-1}(\lambda))=1/\lambda$, as required. ∎
Note that Lemma 2.4 defines the curvature of the hypercircles contained in the thick $\lambda$-sausage defined in Definition 1.3.
3 Proof of Theorems 1.2 and 1.3
The proof of Theorem 1.2 is divided into 2 steps. The strategy is as follows:
@@
    Given a thick $\lambda$-concave body $K$ we construct a convex body $L$ such that $L\subset K$. We do that by using the notion of strict convexity from [11] (Subsection 3.1).
@@
    We use the classical Steiner’s formula to relate the area between $L$ and $K$, and their inner parallel bodies (Subsection 3.2).